OpenStudy (anonymous):

what is the circumference of 17 cm

OpenStudy (camerondoherty):

is 17 cm the diameter or radius?

OpenStudy (anonymous):

Diameter -

OpenStudy (camerondoherty):

Then you multiply \[17 \times \pi\] pi can be represented as 3.14 as a shorter approximation

OpenStudy (anonymous):

Thank you :)

OpenStudy (camerondoherty):

You're Welcome C:

OpenStudy (anonymous):

But how would I begin to work it out fully

OpenStudy (camerondoherty):

You Multiply \[17 \times \pi\]which is also \[17 \times 3.14\]using the approximation

OpenStudy (anonymous):

Oh okay thank you so much

OpenStudy (camerondoherty):

Pleasure to Help c: